OpenStudy (anonymous):

negative 1 over 2 x + 1 = –x + 8

OpenStudy (marihelenh):

\[\frac{ -1 }{ 2x+1 }=-x+8\] Correct?

OpenStudy (marihelenh):

First multiply both side by 2x-1 to get rid of the denominator.